OpenStudy (anonymous):

Sam's plasma TV has a 14 in by 18 in screen with a plastic trim border. The total area of the TV and border is 480 in2. How wide is the trim around the TV?

OpenStudy (anonymous):

(14 + w)(18 + w) = 480 252 + 32w + w^2 = 480 w^2 + 32w - 228 = 0 ( w + 38)(w - 6) = 0 w = 6

OpenStudy (anonymous):

is that correct